# RFUAV: A Benchmark Dataset for Unmanned Aerial Vehicle Detection and Identification

### v1

@@ Appendix 0.E Color Map Detail
In this section, we examine the impact of different color mapping methods and SNR levels on the quality of time-frequency spectrograms. Using the dataset from the experimental subsection, we generate spectrograms for the raw signals of the DJI FPV COMBO drone at SNR levels of 20 dB, 10 dB, 0 dB, -10 dB, and -20 dB. These spectrograms are visualized using 4 types of color map methods: the Autumn method, which colors the STFT results from RGB (255, 0, 0) to RGB (255, 255, 0); the Hot method, which colors the STFT results from RGB (0, 0, 0) to RGB (255, 255, 255); the HSV method, which colors the STFT results from RGB (255, 0, 0) to RGB (255, 0, 255); and the Parula method, which colors the STFT results from RGB (62, 38, 168) to RGB (255, 255, 0). These visualizations are illustrated in Fig. 12.
@@
As shown in the figure, as the SNR decreases, drone signals are gradually overwhelmed by background noise. When the SNR drops to approximately -10 dB, drone signals become almost indiscernible in the spectrogram. Color mapping methods also significantly affect the visual identification of drone signals, as the contrast differences caused by various mapping methods can be substantial. Based on the contrast differences observed in the comparative images, the ranking is as follows: Parula $<$ Autumn $<$ Hot $<$ HSV.
The experimental results shown in Sec. 5.1.3 indicate that color mapping methods with higher contrast differences tend to rely more heavily on abundant training data and exhibit poorer generalization performance.

### v2

@@ Appendix 0.E Color Map Detail
In this section, we examine the impact of different color mapping methods and SNR levels on the quality of time-frequency spectrograms. Using the dataset from the experimental subsection, we generate spectrograms for the raw signals of the DJI FPV COMBO drone at SNR levels of 20 dB, 10 dB, 0 dB, -10 dB, and -20 dB. These spectrograms are visualized using 4 types of color map methods: the ”Autumn” method, which colors the STFT results from RGB (255, 0, 0) to RGB (255, 255, 0); the ”Hot” method, which colors the STFT results from RGB (0, 0, 0) to RGB (255, 255, 255); the ”HSV” method, which colors the STFT results from RGB (255, 0, 0) to RGB (255, 0, 255); and the ”Parula” method, which colors the STFT results from RGB (62, 38, 168) to RGB (255, 255, 0). These visualizations are illustrated in Fig. 12.
@@
As shown in the figure, as the SNR decreases, drone signals are gradually overwhelmed by background noise. When the SNR drops to approximately -10 dB, drone signals become almost indiscernible in the spectrogram. Color mapping methods also significantly affect the visual identification of drone signals, as the contrast differences caused by various mapping methods can be substantial. Based on the contrast differences observed in the comparative images, the ranking is as follows: ”Parula” $<$ ”Autumn” $<$ ”Hot” $<$ ”HSV”.
The experimental results shown in Sec. 5.1 indicate that color mapping methods with higher contrast differences tend to rely more heavily on abundant training data and exhibit poorer generalization performance.